# Solar carbon abundance from 3D non-LTE modelling of the diagnostic lines of the CH molecule

### v1

@@ 6 Conclusions
In this work, we explore the influence of 3D NLTE effects on the formation of diagnostic absorption lines of the CH molecule in the solar spectrum. Our linelist includes the optical CH lines at $4218\text{\,}$-$4356\text{\,}\mathrm{\text{Å}}$ and the far-IR CH lines at 33 $3\,025\text{\,}$-$37\,944\text{\,}\mathrm{\text{Å}}$). Our goal is to test the influence of these effects on the analysis of solar photospheric C abundance. This quantity is particularly relevant in studies of the structure of the solar interior (SerenelliBasu2009; MaggBergemann2022) and in the context of neutrino measurements (AppelBagdasarian2022; BasilicoBellini2023).
We model NLTE radiation transfer in the CH molecule using 3D radiation-hydrodynamics models of the solar atmosphere from our previous work (EitnerBergemann2024) and from the STAGGER grid (MagicCollet2013). We employ an updated version of the CH model molecule from PopaHoppe2023, but grouping individual energy states into superlevels. We have verified carefully that our level averaging procedure does not compromise NLTE effects. For a comparative analysis, we calculate 1D LTE, 1D NLTE, <3D> LTE, 3D LTE and full 3D NLTE CH line profiles. The resulting synthetic spectra are compared with the observed high-resolution solar flux data as well as with the spatially-resolved solar spectra IAG taken at multiple angles across the solar disc (EllwarthSchafer2023).
@@
With this paper, we release the model molecule, our 3D Dispatch model atmosphere is public (EitnerBergemann2024,  [^3] https://nlte.mpia.de) and the line list is provided in tables 1 and 3. Finally we are preparing a dedicated paper to release our updated 3D NLTE spectrum synthesis code.
@@ Acknowledgements
The Spacelab-3 ATMOS atlas ($\mu=0.935$) (FarmerNorton1989) as uploaded by SeoKim2007 (file solspec2b_1.txt) but renormalised using SUPPNet (RozanskiNiemczura2022) will be made available in machine-readable format via CDS upon publication of the paper. The employed model of the CH molecule is uploaded as supplementary material to the paper as well.
@@
Figure 16: IAG observations of optical A-X transition at $4274.19\text{\,}\mathrm{\text{Å}}$ at various $\mu$-angles (EllwarthSchafer2023). The angles starting from $\mu=0.99$ have been offset by 0.1 on the y-axis with respect to one another. The grey observations have been omitted in the final abundance analysis as these are of comparatively poor SNR.
@@
Figure 20: Ratio of molecular number densities for different molecules in an <3D> M3DIS snapshot over the average molecular number density of the full 3D snapshot as a function of optical depth. CH and C2 are the only molecules where using a <3D> atmosphere overestimates their average number density, thereby defying the prediction of UitenbroekCriscuoli2011. See section 4.3 for more details.
@@ A.1 Updates to the CH model molecule
However, the calculations in PH23 were performed using the MULTI code (Carlsson1986), which requires the dissociated "continuum" state to have the highest energy among all states in the model atom/molecule. Due to this limitation, the authors replaced the dissociation threshold of $3.47\text{\,}\mathrm{eV}$ with $5.16\text{\,}\mathrm{eV}$. This incorrect dissociation energy only affected the calculation of collisional dissociation rates. The Seaton1962 and Drawin1969 rates include a factor of $\exp(-E_{ij}/(k_{B}T))$, where $E_{ij}$ is the energy difference between the lower and upper states. Consequently, the collisional dissociation rates were significantly underestimated, leading to an overall overestimation of the NLTE abundance corrections in PH23.
@@ A.5 Infra-red lines: Fitting procedure and error analysis
Figure 23: 1D LTE (MARCS) and 3D NLTE (M3DIS) best-fits of optical diagnostic A-X transitions to the KPNO disk-center (brault1987spectral) and flux (KuruczFurenlid1984) atlases. The legend labels show the corresponding C abundance. The vertical lines frame the windows used in the fitting-procedure.
Figure 24: Spacelab-3 ATMOS atlas observations (FarmerNorton1989) and 3D NLTE (M3DIS) line profile fits of infra-red vibration-rotation lines. The legend labels show the best-fit C abundance. The associated uncertainty is dominated by the uncertainty in the continuum placement. The vertical lines frame the windows used in the fitting-procedure.

### v2

@@ 6 Conclusions
In this work, we explore the influence of 3D NLTE effects on the formation of diagnostic absorption lines of the CH molecule in the solar spectrum. Our linelist includes the optical CH lines at $4218\text{\,}$-$4356\text{\,}\mathrm{\text{Å}}$ and the far-IR CH lines at 33 $3\,025\text{\,}$-$37\,944\text{\,}\mathrm{\text{Å}}$). Our goal is to test the influence of these effects on the analysis of solar photospheric C abundance. This quantity is particularly relevant in studies of the structure of the solar interior (Serenelli et al., 2009; Magg et al., 2022) and in the context of neutrino measurements (Appel et al., 2022; Basilico et al., 2023).
We model NLTE radiation transfer in the CH molecule using 3D radiation-hydrodynamics models of the solar atmosphere from our previous work (Eitner et al., 2024) and from the STAGGER grid (Magic et al., 2013). We employ an updated version of the CH model molecule from Popa et al. (2023), but grouping individual energy states into superlevels. We have verified carefully that our level averaging procedure does not compromise NLTE effects. For a comparative analysis, we calculate 1D LTE, 1D NLTE, <3D> LTE, 3D LTE and full 3D NLTE CH line profiles. The resulting synthetic spectra are compared with the observed high-resolution solar flux data as well as with the spatially-resolved solar spectra IAG taken at multiple angles across the solar disc (Ellwarth et al., 2023).
@@
With this paper, we release the model molecule, our 3D Dispatch model atmosphere is public (Eitner et al., 2024,  [^3] https://nlte.mpia.de) and the line list is provided in tables 1 and 3. Finally we are preparing a dedicated paper to release our updated 3D NLTE spectrum synthesis code.
@@ Acknowledgements
The Spacelab-3 ATMOS atlas ($\mu=0.935$) (Farmer and Norton, 1989) as uploaded by Seo et al. (2007) (file solspec2b_1.txt) but renormalised using SUPPNet (Różański et al., 2022) will be made available in machine-readable format via CDS upon publication of the paper. The employed model of the CH molecule is uploaded as supplementary material to the paper as well.
@@
Figure 16: IAG observations of optical A-X transition at $4274.19\text{\,}\mathrm{\text{Å}}$ at various $\mu$-angles (Ellwarth et al., 2023). The angles starting from $\mu=0.99$ have been offset by 0.1 on the y-axis with respect to one another. The grey observations have been omitted in the final abundance analysis as these are of comparatively poor SNR.
@@
Figure 20: Ratio of molecular number densities for different molecules in an <3D> M3DIS snapshot over the average molecular number density of the full 3D snapshot as a function of optical depth. CH and C2 are the only molecules where using a <3D> atmosphere overestimates their average number density, thereby defying the prediction of Uitenbroek and Criscuoli (2011). See section 4.3 for more details.
@@ A.1 Updates to the CH model molecule
However, the calculations in PH23 were performed using the MULTI code (Carlsson, 1986), which requires the dissociated "continuum" state to have the highest energy among all states in the model atom/molecule. Due to this limitation, the authors replaced the dissociation threshold of $3.47\text{\,}\mathrm{eV}$ with $5.16\text{\,}\mathrm{eV}$. This incorrect dissociation energy only affected the calculation of collisional dissociation rates. The Seaton (1962) and Drawin (1969) rates include a factor of $\exp(-E_{ij}/(k_{B}T))$, where $E_{ij}$ is the energy difference between the lower and upper states. Consequently, the collisional dissociation rates were significantly underestimated, leading to an overall overestimation of the NLTE abundance corrections in PH23.
@@ A.5 Infra-red lines: Fitting procedure and error analysis
Figure 23: 1D LTE (MARCS) and 3D NLTE (M3DIS) best-fits of optical diagnostic A-X transitions to the KPNO disk-center (Brault and Neckel, 1987) and flux (Kurucz et al., 1984) atlases. The legend labels show the corresponding C abundance. The vertical lines frame the windows used in the fitting-procedure.
Figure 24: Spacelab-3 ATMOS atlas observations (Farmer and Norton, 1989) and 3D NLTE (M3DIS) line profile fits of infra-red vibration-rotation lines. The legend labels show the best-fit C abundance. The associated uncertainty is dominated by the uncertainty in the continuum placement. The vertical lines frame the windows used in the fitting-procedure.